Subject: 3.5-m Users Committee minutes 9/15/08
From: Suzanne Hawley
Submitted: Wed, 24 Sep 2008 14:14:23 -0700 (PDT)
Message number: 1112
(previous: 1111,
next: 1113
up: Index)
APO 3.5-M Users Committee Phonecon, 09/15/08
Attending: Suzanne Hawley, Michael Strauss, John Bally, Mark Klaene,
Bill Ketzeback
Absent: Jon Holtzman, Russet McMillan, Bruce Gillespie, Al Harper,
Jon Fulbright, Remy Indebetouw, Rene Walterbos, Scott Anderson
Minutes taken by Bill Ketzeback
**********************************
User feedback, comments from institutional representatives:
Princeton (Strauss) - nothing to report
UW (Hawley) - users are happy with the new TripleSpec data reduction
package which incorporates OH skylines for wavelength calibration. There
is a lot of interest in getting Agile working with TUI to allow easier
observing. Several UW people are interested in MARVELS follow-up of
planet candidates, especially for transit observations - ARC users who
want to collaborate on followup, as for the SDSS-II supernova survey,
should contact Andy Becker and Eric Agol at UW. Florida (Jian Ge)
is interested in bringing a high resolution IR spectrograph and/or an
optical clone (single fiber) of the MARVELS ET instrument to the 3.5m.
Either of these could be used for precision redial velocity followup.
Colorado (Bally) - A Colorado user reported problems with TripleSpec during
the July shared-risk run and was unhappy with the state of the Observing
Specialist training and documentation at that time. Another issue is the
use of the telescope during marginal weather conditions, and getting
feedback from the Observing Specialist as sky conditions are changing.
Also, suggested reviewing procedures for instrument and shift changes on
weekends. Concern that observing programs were unfairly impacted by
scheduling of APOLLO.
*************************************
Discussion of telescope/instruments report - Klaene
The report is given below. Mark provided more information on the
following issues.
The tertiary mirror installation debugging is continuing with work on problems
associated with the axial position and rotation.
Another tropical storm system just hit, bring the total to four over the
summer. The extra rain has reduced the drought but increased the growth of
smaller plants in the forest, which act as fuel sources when they dry out.
So the long term affect of all this moisture may not have lessened the
fire danger as much as might be thought. Prediction for winter is light
snow pack and warmer temperatures.
The DIS red camera has been giving us problems with ion pumps and cryotiger.
We may need to service DIS again soon.
GFP is warm and will stay that way until November.
***********************************************************************
Agile plans for fall 2008 - Klaene
Agile is planned to move to the TR2 port. All parts are on hand, but
we are waiting for the control software to be completed and tested on
the bench. We hope to install it during October engineering time.
Russell Owen is integrating Agile with new ICC and into Hub/TUI, which
hopefully will also be ready by October.
**********************************************************************
Triplespec scheduling for Q4 - Hawley
Triplespec has one more block-scheduled run, with UVa support available,
in October. After that it will be in normal usage on a half-night basis
like the other instruments, and without readily available UVa support.
Users should expect additional time for Tspec instrument changes; these
are reflected in the Q4 schedule which gives 18 minutes between programs
rather than the normal 12. The non-Tspec user is charged 6 minutes
of the instrument change time, as usual, and the Tspec user is charged
12 minutes. We are working on shortening the instrument change time,
but due to the length and weight of the instrument it will likely
always take a few minutes longer.
**********************************************************************
Update on ARCSAT (formerly known as PT) - Klaene, Hawley
The 0.5m telescope formerly known as the Photometric Telescope
(it was used for SDSS and SDSS-II photometry calibration) was formally
turned over to ARC after SDSS II completed operations in July. It
is now known as the ARC Small Aperture Telescope (ARCSAT).
A proposal to support ARCSAT with minimal site operations budget impact
(electricity, internet) and on a low-priority non-interference basis
by existing site staff has been prepared by Mark, Bruce and Suzanne
and will be presented to the ARC BoG in November. Funding for telescope
and instrument improvements, observing, etc. must come from people who
want to use the telescope. Steph Snedden has been appointed ARCSAT Director
and she will be issuing calls for science proposals and scheduling use
of the telescope, pursuant to cash and in-kind contributions. Interested
users from the ARC community should contact Steph and/or Mark and Suzanne.
***********************************
Actions from earlier meetings - group
Science highlights: Please send to Bruce and Suzanne ASAP for the BoG report
ACTION ITEMS:
[open from previous months]: none
[new actions from last month]:
ACTION: Committee members should send one or two science highlights
from the telescope users from their institution to Suzanne
(slh/astro.washington.edu) and Bruce (Gillespie/apo.nmsu.edu)
not later than 1 October for use in the Annual Report to the ARC
Board of Governors meeting in November.
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights
8/1/08 through 9/5/08
Mark Klaene
0) Overview
The summer shutdown was completed one night behind schedule, in part
due to numerous small problems coming out of shutdown, compounded by
uncooperative weather. One half night of DD time was given to engineering.
All shutdown tasks and related projects were completed, but most of the
characterization of the new tertiary mount will be postponed until the
engineering time in September.
The continuing summer rains hindered post-shutdown observing for much
of this period. During this summer, we saw three tropical storms affect
New Mexico's weather.
We have issued a contract for a new heating boiler for Dorm B, and repair
of the fire alarm system, which was damaged by lightning.
1) Telescope
The Primary, Secondary, and Tertiary mirrors were washed. We installed
new tertiary mounts and actuators for piston and tip/tilt. This new
design eliminates a high-rate failure item, the flex pivots. It also i
ncludes more and better mirror safeties, stiffer support, and
a more robust control system.
2) Instruments
DIS and SPIcam were vacuum-pumped and a new ion pump installed on DIS
Red camera. The Echelle, TripleSpec, and NIC-FPS instruments remained cold
and powered-on during the shutdown. When time becomes available to refine
the control logic, Agile will be moved to the TR2 port (hopefully in
October). GFP is currently warm (no planned near-term use), and APOLLO
saw a servicing trip by the PI team and remains operational.
3) CIF projects
The tertiary support/actuator upgrade is installed and operating, with
additional characterization planned. We have been able to lower the
tertiary rotation limit from above 80 degrees to above 35 degrees altitude.
The problem below 35 degrees is with the rotation axis mechanicals;
we have a potential fix for that that needs implementation and testing.
Being able to rotate the tertiary to different ports over a wide range
of telescope altitudes will help improve observing efficiency.
***********************************
Next meeting:
The next Users Committee phonecon will be on Monday, 6 October, at 8:30 AM
Pacific Time. The agenda and other materials will be sent to the committee
members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1112 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO